OpenStudy (anonymous):

Find the quantity if v = 5i - 7j and w = 3i + 2j. a) ||v||+||w|| b) ||v||-||w||

OpenStudy (anonymous):

We wish to find the value of ||v+w|| if v=5i-7j and w=3i+2j ........_______ |r| =../r.+..r .....√...1....2 if r = r(1)i +r(2)j We first find the value of v+ w by adding v(1) +w(1) = 5+3 = 8 and by adding v(2) +w(2) = -7+2 = -5 so v+w = 8i -5j and .............______ |v+w| = √8²+(-5)² ...._______ =√64 + 25 ....___ =√89

OpenStudy (anonymous):

||v|| = \[\sqrt{5^{2}+(-7^{2}})\] ||w|| = \[\sqrt{3^{2}+2^{2}}\]

OpenStudy (anonymous):

no no no \[||v+w||\neq||v||+||w||\]

OpenStudy (anonymous):

\[||v|| = \sqrt{74}\] \[||w||=\sqrt{13}\]

OpenStudy (anonymous):

so does that mean that \[||v||+||w|| = \sqrt{74}+\sqrt{13}\]

OpenStudy (anonymous):

and \[||v||-||w|| = \sqrt{74}-\sqrt{13}\]

OpenStudy (tkhunny):

That's it. Good work. Why do you doubt?

OpenStudy (anonymous):

well, I'm not sure if I should or can simplify further, is that answer acceptable?

OpenStudy (tkhunny):

Fine. Not anywhere else to go with it.

OpenStudy (anonymous):

perfect! Thank you!